Suppose a random selected family has 3 children. Define X to be the number of boys in the family, X = 0,1,2, and 3. Write the sample space, construct a probability distribution in table form, and make histogram.

1
Expert's answer
2022-03-19T02:40:13-0400

Sample space: X = {0,1,2,3}

X ~ Bin(3, 0.5)

"P(X=0)={3 \choose 0}*({\frac 1 2})^3={\frac 1 8}"

"P(X=1)={3 \choose 1}*({\frac 1 2})^3={\frac 3 8}"

"P(X=0)={3 \choose 2}*({\frac 1 2})^3={\frac 3 8}"

"P(X=0)={3 \choose 0}*({\frac 1 2})^3={\frac 1 8}"